If I traveled 60 mph and was an hour early, and traveled 30 mph and was an hour late how far was the airport?
Answer by jim_thompson5910(35256)

Let t = time it takes to travel a distance of d miles (and be on time)

So to be an hour early, it would mean that you travel 1 hour less than normal, which means that you spend t-1 hours on the road.

To be an hour late, you add 1 to t to get t+1 hours.

-------------------------------------------------------

Being 1 hr early

d = rt

d = 60t

d = 60(t-1)

-----------------------------

Being 1 hr late

d = rt

d = 30t

d = 30(t+1)

---------------------------

Equate the two and solve for t


60(t-1) = 30(t+1)

60t-60 = 30t+30

60t-30t = 30+60

30t = 90

t = 90/30

t = 3

So if you arrive on time, then you will spend 3 hours on the road. Being early means you only spend 2 hours on the road and being late means you spend 4 hours on the road.

Use this to find d

d = 60(t-1)

d = 60(3-1)

d = 120

Therefore, the distance is 120 miles.